# Characterizing Multimodal Long-form Summarization: A Case Study on Financial Reports

### v1

@@ Abstract
As large language models (LLMs) expand the power of natural language processing to handle long inputs, rigorous and systematic analyses are necessary to understand their abilities and behavior. A salient application is summarization, due to its ubiquity and controversy (e.g., researchers have declared the death of summarization). In this paper, we use financial report summarization as a case study because financial reports not only are long but also use numbers and tables extensively. We propose a computational framework for characterizing multimodal long-form summarization and investigate the behavior of Claude 2.0/2.1, GPT-4/3.5, and Command. We find that GPT-3.5 and Command fail to perform this summarization task meaningfully. For Claude 2 and GPT-4, we analyze the extractiveness of the summary and identify a position bias in LLMs. This position bias disappears after shuffling the input for Claude, which suggests that Claude has the ability to recognize important information. We also conduct a comprehensive investigation on the use of numeric data in LLM-generated summaries and offer a taxonomy of numeric hallucination. We employ prompt engineering to improve GPT-4’s use of numbers with limited success. Overall, our analyses highlight the strong capability of Claude 2 in handling long multimodal inputs compared to GPT-4.
@@ 1 Introduction
Second, given the important role of numbers, we conduct a thorough investigation on the use of numbers in LLM-generated summaries. We find that small models tend to include generic texts without mentioning any specific numbers. In comparison, large models use numbers more extensively and often perform simple operations on numbers such as rounding and taking the difference. Claude 2 outperforms GPT-4 with a higher tabular numbers utilization ratio and a larger number density. We further provide a taxonomy of numeric hallucinations through manual annotation. The results show that although LLMs hallucinate about 5% of the time, they may fail to capture the semantic relationship between numeric data and textual data, an instance of context mismatch.
Finally, we explore the promise of prompt engineering to enhance the use of numbers by GPT-4. We find that GPT-4 extracts more numbers when prompted properly, but still uses fewer tabular numbers than the summary generated by a simple prompt with Claude 2.
@@
  • We develop a computational framework for characterizing multimodal long-form summarization that accounts for both numeric values and sentence coverage.
  • We demonstrate that Cohere and GPT-3.5 cannot perform such long-form summarization.
  • We compare the behavior and Claude and GPT-4, and show that Claude demonstrates a stronger ability to use numbers and is less extractive.
@@ 2 Dataset and Methods
We randomly selected 1,000 HTML files of 10-K forms obtained from the Electronic Data Gathering, Analysis, and Retrieval (EDGAR) system. These files are then converted into clean JSON format using EDGAR-CRAWLER (Loukas et al., 2021) and we use Item 7 as the target report to do summarization. Additionally, to explore whether LLMs have position biases, we randomly shuffle the paragraphs of each report treating tables as single paragraphs, and create a new shuffled version for each report.
Models. We evaluate summaries generated by five commercial models with top context window length (wl) in July 2023 when we started this work: (1) Claude 2.0 (claude-2.0, wl=100K tokens), (2) Claude 2.1 (claude-2.1, wl=200K tokens), [^1] We added Claude 2.1 when it was out, but Claude 3 came out in February 2024, so we have not been able to include it. (3) GPT-3.5 (gpt-3.5-turbo-1106, wl=16K tokens), (4) GPT-4 (gpt-4-1106-preview, wl=128K tokens) (Achiam et al., 2023), and (5) Cohere (command, wl=100K characters). For Claude 2.0 and Claude 2.1, we set max output tokens to 4,096 and temperature to 1. GPT-3.5 and GPT-4 are used with default hyperparameters. We use the Cohere co.summarize API and set the output summary length to “long”, format as bullets, high extractiveness, and temperature to 0.3 according to the official recommendation. Except for Cohere, which does not need additional prompts with the specific endpoint, we use simple prompts for summarization. Example prompts are shown in Appendix D.
General analysis of summary. Table 1 presents basic statistics of the report and summary generated by each model. Notably, the summaries produced by GPT-4 exhibit the longest average word length of 421.51. Claude 2.1, with its substantial 200K-token context length, extracts the highest number of numeric values. A more detailed investigation of numeric values utilization will be presented in Section 4. It is worth mentioning that certain reports require truncation to meet the context limit of different models, with all reports truncated for GPT-3.5, 95.9% for Cohere, 4.1% for Claude 2.0, 0.6% for GPT-4, and no truncation for Claude 2.1.
@@ 3.1 To What Extent Are Generated Summaries Extractive?
Type
Sentences
Score
1-1
summary sentence: As of March 31, 2020 , accumulated deficit was \$112.3 million .
1.70
@@
summary sentence: - For 2019 , capital expenditures are estimated at \$2.9 billion .
0.95
@@
2-1
summary sentence: - Operating cash flows increased by \$98 million due to higher net working capital inflows .
0.94
@@
summary sentence:- Investing cash flows decreased due to lower cash paid for acquisitions .
0.92
@@
Abstractive
summary sentence: Results of operations : A detailed comparison between the years 2018 and 2017 displays an increase in net income and net sales , while gross profit as a percentage of net sales experienced a decrease.
N/A
@@
We modify the coverage measure in Grusky et al. (2018) to measure the similarity between two sentences with a greedy match algorithm. Given a summary sentence $S=\{t_{S1},t_{S2},...,t_{Sn}\}$ and a report sentence $R=\{t_{R1},t_{R2},...,t_{Rm}\}$ after removing stopwords, where $t_{i}$ is a token, the length of $S$ is $n$, and the length of $R$ is $m$. At a high level, $\operatorname{similarity}(S,R)$ is defined as the ratio of matched tokens with a quadratic bonus that rewards longer-matched sequences. Specifically, for each summary sentence, this algorithm processes its tokens iteratively, discovering and aligning the longest matching token sequences from each report sentence. Then, for each report sentence $R$, we get a list of the longest matching token sequences $M(S,R)=\{\{t_{i},t_{i+1},...,t_{i+p}\},...,\{t_{j},t_{j},...,t_{j+q}\}|t\in S\cap R\}$ compared with the summary sentence $S$. The similarity score is computed as:
@@
Following a meticulous manual check (Appendix F), we find sentences with a similarity score above 0.8 to contain almost verbatim excerpts from the source text, which we define as extractive. We calculate the similarity score between each summary sentence and report sentence. If the top similarity score is greater than 0.8, we consider the summary sentence a 1-1 extractive sentence. For the rest of the summary sentences, we compute the similarity score with a combination of two report sentences (2-1) and call it a 2-1 synthesizing sentence if it has a similarity score larger than 0.8. Examples of 1-1 extractive sentences and 2-1 synthesizing sentences are shown in Table 2.
Extractive sentences represent 30% to 40% of the summary. Table 3 shows the percentage of extractive sentences. Claude 2.1 generates the most extractive contents with 41.50% of extractive summary sentences, which is much higher than 31.24% of Claude 2.0. This may be a result of the multi-level bullet formats of Claude 2.1. GPT-4 generates comparable extractive contents with Claude 2.0 but with a smaller percentage of 2-1 synthesizing sentences. Summaries generated by Claude 2 for the shuffled report exhibit more extractiveness, with 6-7% higher proportions of extractive sentences compared with that of original reports. In contrast, GPT-4 generates fewer extractive sentences from shuffled reports. This observation demonstrates differing behavior of Claude and GPT-4 when processing incoherent texts.
Analysis of the remaining abstractive sentences. To gain insights into the abstractive capabilities of LLMs, we conduct a case study for the abstractive sentences present in the summaries. Specifically, we find that the GPT-4’s summary shows a high degree of condensation such as summarizing net income, net sales, and gross profit in a single sentence (Table  2). It paraphrases the original information across various locations and compacts it into one single sentence, generating new phrases such as “display an increase in net income”. The entire meaning is retained, but the content organization is significantly different from the original report. See the Appendix G for similar examples from Claude 2.
@@ 3.2 Where Does the Information Come From?
Identify source contents. Based on the extractive analysis, we analyze the distribution of the 1-1 extractive sentences. For each summary sentence, the report sentence with the highest similarity scores is treated as its source. If there exist several report sentences with the same score, we choose the one with the top cosine-similarity using sentence embeddings generated by SentenceTransformers (Reimers & Gurevych, 2019). Then we calculate the position of each report sentence and visualize the distribution in Figure 1. We also run the same analysis by including the location of 2-1 synthesizing summary sentence sources. The results are similar and presented in Appendix B.
@@
Shuffled report analysis. To understand whether the observed pattern is a bias in LLM’s behavior or a consequence of document structure (there is an overview at the beginning of Item 7), we analyze the summaries after we shuffle the input reports. Figure 1(b) shows that the source information is evenly distributed for Claude 2. The distribution becomes heavy in the beginning when comparing the summary of the shuffled report with the original report, suggesting that Claude 2 seems to recognize the important information even after shuffling. In comparison, GPT-4 maintains a strong position bias for the shuffled input: the summary of the shuffled input now favors the beginning part despite that it is essentially arbitrary text from the input. [^2] Due to cost reasons, the GPT-4 analysis is based on 100 shuffled reports.
@@ 4.1 Basic Numeric Values Analysis
Claude 2 demonstrates a more sophisticated use of numbers than GPT-4. As illustrated in Table 1, Claude 2.1 demonstrates a strong ability to use tabular numeric values, with 8.37% of its numeric values belonging to type B, compared to only 4.98% for GPT-4. The difference is statistically significant according to a two-sample $t$-test ($p<0.001$). However, it is noteworthy that in the reports, more than half of the numeric values were exclusively present in tabular formats, suggesting that LLMs tend to prioritize numeric values found in the textual narratives and pay relatively less attention to tabular data.
@@
Figure 2: The summary sentence presents a rounded figure with the unit adjusted as per the remark preceding the table.
Type D numbers can come from numeric operations. Numeric values of type D, which represent mismatches between the summary and the report, accounted for an average of 2.76 instances per summary generated by Claude 2.1 and 2.66 for Claude 2.0. In contrast, GPT-4 has only 0.46 D numbers. Through human annotation, we identified that these mismatches often stem from simple operations such as rounding, calculating differences, or computing rates of change (e.g., 86.36% of type D numbers generated by Claude 2.0 come from such operations).
@@ 4.2 Numeric Hallucinations
  • Arithmetic Error: Numbers generated from incorrect mathematical calculations applied to numeric values from the source text.
  • Context mismatch: An inconsistency where the same numeric value applies to different contexts in the source vs. in the summary.
@@ 5 Prompt Engineering to Improve Use of Numbers
Model Prompt Words TotalNum Percentage %
@@ A B C D
GPT-4 simple 421.51 5.81 57.47 4.91 29.77 7.85
NUM 423.50 14.30 56.64 4.90 24.48 13.99
TAB 429.70 9.40 62.77 2.13 25.53 9.57
NUMTAB 421.51 18.90 55.03 6.88 21.16 16.93
Claude 2.0 simple 223.82 11.09 41.64 8.27 26.10 23.99
Claude 2.1 simple 229.66 11.56 40.81 8.37 26.91 23.88
@@
In order to enhance GPT-4’s performance in extracting numeric values, particularly from tables, we design several prompts: NUM to explicitly request the inclusion of numeric values. TAB to explicitly request the inclusion of tabular numbers. NUMTAB to explicitly request both numeric values and tabular numbers. See Appendix D for details of these prompts.
Table 5 shows that the NUMTAB prompt enables GPT-4 to extract the highest number of numeric values and tabular numbers (type B) across all prompt strategies. However, when prompted to include tabular numbers (TAB prompt), GPT-4 still prioritizes the extraction of type A numbers, as evidenced by the 62.77% of type A numbers. Consequently, the percentage of type B numbers extracted is always lower compared to the summary generated by Claude 2.1 using a simple prompt. Overall, it seems that Claude still demonstrates a stronger ability to use numbers in summarization.
@@ 6 Related Work
7 Conclusion
We present the first study to characterize summaries generated by LLMs for multimodal long-form inputs. Our results reveal that different LLMs may approach this task quite differently in terms of extractiveness, position bias, and use of numbers. While the position bias of GPT-4 is probably sub-optimal, it is unclear what makes a good summary, especially for such long-form inputs. Our community urgently need novel perspectives of evaluating summarization. In other words, the death of summarization might be the death of simple relevance and human preference. On the common concern about hallucinations, we find that hallucinations do not happen very frequently. Future work may not only focus on the frequency but also the potential impact of such misuse.

### v3

@@ Abstract
As large language models (LLMs) expand the power of natural language processing to handle long inputs, rigorous and systematic analyses are necessary to understand their abilities and behavior. A salient application is summarization, due to its ubiquity and controversy (e.g., researchers have declared the death of summarization). In this paper, we use financial report summarization as a case study because financial reports are not only long but also use numbers and tables extensively. We propose a computational framework for characterizing multimodal long-form summarization and investigate the behavior of Claude 2.0/2.1, GPT-4/3.5, and Cohere. We find that GPT-3.5 and Cohere fail to perform this summarization task meaningfully. For Claude 2 and GPT-4, we analyze the extractiveness of the summary and identify a position bias in LLMs. This position bias disappears after shuffling the input for Claude, which suggests that Claude seems to recognize important information. We also conduct a comprehensive investigation on the use of numeric data in LLM-generated summaries and offer a taxonomy of numeric hallucination. We employ prompt engineering to improve GPT-4’s use of numbers with limited success. Overall, our analyses highlight the strong capability of Claude 2 in handling long multimodal inputs compared to GPT-4. The generated summaries and evaluation code are available at https://github.com/ChicagoHAI/characterizing-multimodal-long-form-summarization.
@@ 1 Introduction
Second, given the important role of numbers, we conduct a thorough investigation on the use of numbers in LLM-generated summaries. We find that small models like GPT-3.5 tend to include generic texts without mentioning any specific numbers. In comparison, large models use numbers more extensively and often perform simple operations on numbers such as rounding and taking the difference. Claude 2 outperforms GPT-4 with a higher tabular numbers utilization ratio and a larger number density. We further provide a taxonomy of numeric hallucinations through manual annotation. The results show that although LLMs hallucinate about 5% of the time, they may fail to capture the semantic relationship between numeric data and textual data, an instance of context mismatch.
Finally, we explore the promise of prompt engineering to enhance the use of numbers by GPT-4. We find that GPT-4 extracts more numbers when prompted properly, but still uses fewer tabular numbers than the summary generated by a simple prompt with Claude 2. Also, fewer hallucinations are generated when using Chain-of-Thought prompting by GPT-4.
@@
  • We develop a computational framework for characterizing multimodal long-form summarization that accounts for both information usage and numeric values.
  • We demonstrate that Cohere and GPT-3.5 cannot perform such long-form multimodal summarization.
  • We compare the behavior of Claude and GPT-4, and show that Claude demonstrates a stronger ability to use numbers and seems to recognize important information.
@@ 2 Dataset and Methods
We randomly selected 1,000 HTML files of 10-K forms obtained from the Electronic Data Gathering, Analysis, and Retrieval (EDGAR) system. These files are then converted into clean JSON format using EDGAR-CRAWLER (Loukas et al., 2021) and we use Item 7 as the target report to do summarization. Additionally, to explore whether LLMs have position biases, we generate shuffled versions of each report by randomly reordering its paragraphs, treating tables as individual paragraphs.
Models. We evaluate summaries generated by five commercial models with top context window length (wl) in July 2023 when we started this work: (1) Claude 2.0 (claude-2.0, wl=100K tokens), (2) Claude 2.1 (claude-2.1, wl=200K tokens), [^1] We added Claude 2.1 when it was out, but Claude 3 came out in February 2024, so we have not been able to include it. (3) GPT-3.5 (gpt-3.5-turbo-1106, wl=16K tokens), (4) GPT-4 (gpt-4-1106-preview, wl=128K tokens) (Achiam et al., 2023), and (5) Cohere (command, wl=100K characters). For Claude 2.0 and Claude 2.1, we set max output tokens to 4,096 and temperature to 1. GPT-3.5 and GPT-4 are used with default hyperparameters. We use the Cohere co.summarize API and set the output summary length to “long”, format as bullets, high extractiveness, and temperature to 0.3 according to the official recommendation. Except for Cohere, which does not need additional prompts with the specific endpoint, we use simple prompts for summarization. Detailed prompts are shown in Appendix D.
General analysis of summary. Table 1 presents basic statistics of the report and summary generated by each model. Notably, the summaries produced by GPT-4 exhibit the longest average word length of 421.51. With its substantial 200K-token context length, Claude 2.1 outperforms other models by extracting an average of 11.56 numeric values per summary. A more detailed investigation of numeric values utilization will be presented in Section 4. It is worth mentioning that certain reports require truncation to meet the context limit of different models, with an averaged of 436, 143, 12,656, and 7,690 words truncated for Claude 2.0, GPT-4, GPT-3.5, and Cohere respectively.
@@ 3.1 To What Extent Are Generated Summaries Extractive?
Type Sentences Score
1-1 summary sentence: As of March 31, 2020 , accumulated deficit was \$112.3 million . 1.70
@@
summary sentence: - For 2019 , capital expenditures are estimated at \$2.9 billion . 0.95
@@
2-1 summary sentence: - Operating cash flows increased by \$98 million due to higher net working capital inflows . 0.94
@@
summary sentence:- Investing cash flows decreased due to lower cash paid for acquisitions . 0.92
@@
Abstractive summary sentence: Results of operations : A detailed comparison between the years 2018 and 2017 displays an increase in net income and net sales , while gross profit as a percentage of net sales experienced a decrease. N/A
@@
We modify the coverage measure in Grusky et al. (2018) to measure the similarity between two sentences with a greedy match algorithm. Given a summary sentence $S=\{t_{S1},t_{S2},...,t_{Sn}\}$ and a report sentence $R=\{t_{R1},t_{R2},...,t_{Rm}\}$ after removing stopwords, where $t_{i}$ is a token, the length of $S$ is $n$, and the length of $R$ is $m$. At a high level, $\operatorname{similarity}(S,R)$ is defined as the ratio of matched tokens with a quadratic bonus that rewards longer-matched sequences. Specifically, for each summary sentence, this algorithm processes its tokens iteratively, discovering and aligning the longest matching token sequences from each report sentence. Then, for each report sentence $R$, we get a list $M(S,R)=\{\{t_{i},t_{i+1},...,t_{i+p}\},...,\{t_{j},t_{j},...,t_{j+q}\}|t\in S\cap R\}$ consisted of the longest matching token sequences $m$ compared with the summary sentence $S$. The similarity score is computed as:
@@
Following a meticulous manual check (Appendix F), we find sentences with a similarity score above 0.8 to contain almost verbatim excerpts from the source text, which we define as extractive. We calculate the similarity score between each summary sentence and report sentence. If the top similarity score is greater than 0.8, we consider the summary sentence a 1-1 extractive sentence. For each remaining summary sentence, we calculate a similarity score with a combination of two report sentences. If this score exceeds 0.8, we call it a 2-1 synthesizing sentence. Examples of 1-1 extractive sentences and 2-1 synthesizing sentences are shown in Table 2.
Extractive sentences represent 30% to 40% of the summary. Table 3 shows the percentage of extractive sentences. Claude 2.1 generates the most extractive contents with 41.50% of extractive summary sentences, which is much higher than 31.24% of Claude 2.0. This may be a result of the multi-level bullet formats of Claude 2.1. GPT-4 generates comparable extractive contents with Claude 2.0 but with a smaller percentage of 2-1 synthesizing sentences. Summaries generated by Claude 2 for the shuffled report exhibit more extractiveness, with 6-7% higher proportions of extractive sentences compared with that of original reports. In contrast, GPT-4 generates fewer extractive sentences from shuffled reports. This observation demonstrates the differing behavior of Claude and GPT-4 when processing incoherent texts.
Analysis of the remaining abstractive sentences. To gain insights into the abstractive capabilities of LLMs, we conduct a case study for the abstractive sentences present in the summaries. Specifically, we find GPT-4 generates abstractive summary sentences with a high degree of condensation, such as summarizing net income, net sales, and gross profit in a single sentence (Table 2). It paraphrases the original information across various locations and compacts it into one single sentence, generating new phrases such as “display an increase in net income”. The entire meaning is retained, but the content organization is significantly different from the original report. See the Appendix G for similar examples from Claude 2.
@@ 3.2 Where Does the Information Come From?
Identify source contents. Based on the extractive analysis, we analyze the distribution of the 1-1 extractive sentences. For each summary sentence, the report sentence with the highest similarity score is treated as its source. If there exist several report sentences with the same score, we choose the one with the top cosine-similarity using sentence embeddings generated by SentenceTransformers (Reimers & Gurevych, 2019). Then we calculate the position of each source report sentence and visualize the distribution in Figure 1. We also run the same analysis by including the location of 2-1 synthesizing summary sentence sources. The results are similar and presented in Appendix B.
@@
Shuffled report analysis. To understand whether the observed pattern is a bias in LLM’s behavior or a consequence of document structure (there is an overview at the beginning of Item 7), we analyze the summaries after we shuffle the input reports. Figure 1(b) SS-SR shows that the source information is evenly distributed for Claude 2. The distribution SS-OR becomes heavy in the beginning when comparing the summary of the shuffled report with the original report, suggesting that Claude 2 seems to recognize the important information even after shuffling. In comparison, GPT-4 maintains a strong position bias for the shuffled input: the summary of the shuffled input now favors the beginning part despite that it is essentially arbitrary text from the input. [^2] Due to cost reasons, the GPT-4 analysis is based on 100 shuffled reports.
@@ 4.1 Basic Numeric Values Analysis
Claude 2 demonstrates a more sophisticated use of numbers than GPT-4. As illustrated in Table 1, Claude 2.1 demonstrates a strong ability to use tabular numeric values, with 8.37% of its numeric values belonging to type B, compared to only 4.98% for GPT-4. The difference is statistically significant according to a two-sample $t$-test ($p<0.001$). However, it is noteworthy that in the reports, more than half of the numeric values were exclusively present in tabular formats, suggesting that LLMs tend to prioritize numeric values found in the text and pay relatively less attention to tabular data.
@@
Figure 2: The summary sentence presents a rounded figure with the unit adjusted as per the remark preceding the table. The input tables are in HTML format as shown in the figure’s bottom section.
Type D numbers can come from numeric operations. Numeric values of type D, which represent mismatches between the summary and the report, accounted for an average of 2.76 instances per summary generated by Claude 2.1 and 2.66 for Claude 2.0. In contrast, GPT-4 has only 0.46 type D numbers. Through human annotation, we identified that these mismatches often stem from simple operations such as rounding, calculating differences, or computing rates of change (e.g., 86.36% of type D numbers generated by Claude 2.0 come from such operations).
@@ 4.2 Numeric Hallucinations
  • Arithmetic Error: Numbers generated from incorrect mathematical calculations applied to numeric values from the report.
  • Context mismatch: An inconsistency where the same numeric value applies to different contexts in the report vs. in the summary.
@@ 5 Prompt Engineering to Improve Use of Numbers
Model Prompt Total Dens.Percentage %
@@ A B C D
GPT-4 simple 5.81 1.38 57.47 4.91 29.77 7.85
NUM 14.30 3.38 56.64 4.90 24.48 13.99
TAB 9.40 2.19 62.77 2.13 25.53 9.57
NUMTAB 18.90 4.48 55.03 6.88 21.16 16.93
CoT 11.30 3.49 43.36 19.47 28.32 8.85
Claude 2.1 simple 11.56 5.03 40.81 8.37 26.91 23.88
@@
To enhance GPT-4’s performance in extracting numeric values, particularly from tables, we design three explicit prompts and one chain-of-thought (CoT) prompt (Appendix D): NUM to explicitly request the inclusion of numeric values. TAB to explicitly request the inclusion of tabular numbers. NUMTAB to explicitly request both numeric values and tabular numbers. CoT to give intermediate reasoning steps of using numeric values.
Figure 3: Hallucinated numbers frequency of summaries generated by simple prompt and CoT prompt respectively.
As shown in Table 5, the NUMTAB prompting strategy enables GPT-4 to extract the highest number of numeric values and achieve the highest summary number density of 4.48% among the five prompt strategies. However, it still falls short of the 5.03% density achieved by Claude 2.1 using a simple prompt. Notably, the CoT prompting substantially enhances GPT-4 using tabular numbers, with 19.47% of the summary numbers being extracted exclusively from the report tables.
To assess the effectiveness of CoT prompting in reducing hallucinations, we annotated a random sample of 10 summaries generated by both GPT-4 and Claude 2.1. As shown in Figure 3, much fewer hallucinations are generated when using CoT prompts to guide GPT-4 on avoiding specific numeric hallucinations, but this is not the case for Claude 2.1. This shows GPT-4’s better ability to leverage CoT prompting for introspecting its own output.
@@ 6 Related Work
7 Conclusion and Discussion
We present the first study to characterize summaries generated by LLMs for multimodal long-form inputs. Our results reveal that different LLMs may approach this task quite differently in terms of extractiveness, position bias, and use of numbers. While the position bias of GPT-4 is probably sub-optimal, it is unclear what makes a good summary, especially for such long-form inputs. On the common concern about hallucinations, we find that hallucinations do not happen very frequently. Future work may not only focus on the frequency but also the potential impact of such misuse. The underlying causes of our observations also remain unclear; we hypothesize that Claude may have more exposure to enterprise data/applications involving structured data and business metrics. This exposure could potentially explain Claude’s stronger ability to use numbers in the summary. However, it is important to note that all the LLMs used in our work are closed and the hypotheses need further investigation.
Our research community urgently need novel perspectives for evaluating summarization. The “death” of summarization is an opportunity for advancing summarization beyond simple relevance and human preference.
Acknowledgements
We thank insightful comments from anonymous reviewers and members of the Chicago Human+AI Lab, especially Chao-Chun Hsu for his early contribution in building the dataset. We sincerely thank Kaiqiao Han, Haoran Deng, and Yang Yang for their valuable advice on our work. This work has been supported in part by J.P. Morgan AI Faculty Research Award. This paper was prepared for informational purposes in part by the Artificial Intelligence Research group of JPMorgan Chase & Co and its affiliates (“J.P. Morgan”) and is not a product of the Research Department of J.P. Morgan. J.P. Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.